OpenStudy (anonymous):

Miguel takes classes at both Westside Community College and Pinewood Community College. At Westside, class fees are $98 per credit hour, and at Pinewood, class fees are $115 per credit hour. Miguel is taking a combined total of 17 credit hours at the two schools. Suppose that he is taking w credit hours at Westside. Write an expression for the combined total dollar amount he paid for his class fees.

OpenStudy (anonymous):

17 hours = westside fees = 98 per credit hour / w pinewood fees = 115 per credit hour 98w + 115 * this is what I have so far not sure If i'm on the right track or what to do next exactly though

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

Does that make sense so far @Cassandra_Lea_96 ?

OpenStudy (anonymous):

yes, thank you =)

OpenStudy (anonymous):

how do I continue this though to find the total he paid

OpenStudy (skullpatrol):

What does the question ask for?

OpenStudy (skullpatrol):

an expression for the combined total dollar amount he paid for his class fees.

OpenStudy (skullpatrol):

Amount paid at West + Amount paid at Pine = combined total

OpenStudy (anonymous):

98w = at westside and 17 - w = at pinewood so 98w + (17 - w) = total?

OpenStudy (skullpatrol):

almost

OpenStudy (skullpatrol):

you missed something

OpenStudy (anonymous):

115 for the pinewood fees?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

Where does that plug into I just wrote it out but pretty sure it wasn't right =O

OpenStudy (anonymous):

It would go in the second part of the equation with the w though right

OpenStudy (skullpatrol):

Amount paid at West + Amount paid at Pine = combined total 98w + 115(17 - w) = total

OpenStudy (anonymous):

Ok that was my next option lol sorry I made that so complicated, but thank you so much for helping me!!

OpenStudy (skullpatrol):

Thanks for trying to learn :)

OpenStudy (anonymous):

Math is NOT my strong subject haha

OpenStudy (skullpatrol):

It takes a lot of practice...

OpenStudy (skullpatrol):

...teachers ALWAYS say that, but you just have to keep trying

OpenStudy (anonymous):

Yeahh give me science, english, writing, psychology of any sort, and I'm good but History If I don't find it interesting is hard and Spanish is ehh but MATH I've alwaaays struggled with

OpenStudy (skullpatrol):

Interest plays a HUGE part in any type of learning :)

OpenStudy (skullpatrol):

You will not remember what you're not interested in...

OpenStudy (skullpatrol):

...just try to fool yourself into being interested at the beginning. It does get really interesting later on.

OpenStudy (anonymous):

I'll keep that in mind =) thank you so much again for being so patient with me!

OpenStudy (skullpatrol):

You did most of the hard work on your own, which is a GREAT sign :)

OpenStudy (anonymous):

thank you feels good to hear haha =)

OpenStudy (skullpatrol):

cya